# Towards Efficient Agents: A Co-Design of Inference Architecture and System

### v1

@@ 1 Introduction
Autonomous agents [chen2025pangu, huang2025deep], powered by large language models (LLMs), are poised to revolutionize domains ranging from deep scientific research and iterative software engineering to knowledge-intensive decision-making. Unlike conventional single-turn LLM applications—where the model generates a one-off response and terminates—autonomous agents operate within a persistent Think–Act–Observe loop. In this cycle, the agent formulates plans, interacts with external tools (e.g., search engines, simulators, code interpreters), observes outcomes, and iteratively reasons based on accumulated context. While this continuous, reflection-driven process endows agents with remarkable problem-solving flexibility, it imposes severe latency penalties and computational overheads. Each “Think” step typically triggers a large model inference involving extended contexts, parallel requests, and complex reasoning paths, resulting in cumulative delays that hinder real-time deployment and scalability.
A fundamental dichotomy exists between agent acceleration and traditional LLM inference optimization. Conventional speedups—such as quantization [li2025kvtuner], pruning [chen2025pangu, pei2025cmoe, pei2024fusegpt], or speculative decoding [sun2025scaling]—predominantly focus on intra-step efficiency, aiming to maximize tokens per second (TPS). In contrast, the efficiency of an autonomous agent is defined by the inter-step time-to-completion of multi-turn tasks. A marginal gain in decoding speed is negligible if the agent, due to degraded reasoning quality, requires dozens of additional correction cycles. Thus, optimizing an agent system demands a paradigm shift: bridging intra-step acceleration (faster inference) with inter-step efficacy (fewer steps to solution) [wu2025resum, lin2025trimr].
Consequently, traditional metrics like TPS can be perilously misleading in agentic contexts. An agent is not merely a passive text generator but an active reasoning system where progress is path-dependent [java2025characterizing]. A “faster but less accurate” model often regresses overall performance by introducing factual errors, invalid tool calls, or incoherent summaries that cascade through subsequent turns. Such errors necessitate retries and redundant inference, negating any low-level speedups. Therefore, agent performance must be evaluated through holistic metrics that consider End-to-End Task Latency (wall-clock time), Total Computational Cost (aggregate FLOPs), and Turn Efficiency (reasoning stability).
@@
We argue that accelerating agent inference requires a synergistic, multi-layered framework spanning both the reasoning process and the inference engine. To this end, we introduce AgentInfer, a hierarchical framework that systematically enhances agent efficiency. While our principles apply to general agentic systems (e.g., coding or dialogue agents), we validate them within the demanding context of Deep Research Agents. At the reasoning level, we propose AgentCollab, a self-evaluation-driven dual-model mechanism. It delegates routine inference and tool execution to a smaller model, while a larger model handles strategic planning and rescues stalled trajectories. This is governed by structured Progress Check signals, which trigger escalation only when necessary, thereby conserving expensive capacity. Complementarily, AgentCompress implements lightweight context summarization, pruning redundant reasoning traces to maintain compact yet semantically complete histories. At the inference-engine level, AgentSAM advances speculative decoding [hu2024samdecodingspeculativedecoding] by using a suffix automaton to efficiently reuse token-level predictions across agent sessions. Simultaneously, AgentSched introduces a hybrid, KVCache-aware scheduling policy. By utilizing a dynamic shadow-price controller, it adaptively switches between Shortest-Job-First and cache-preserving modes, balancing latency and cache hit rates under dynamic workloads. Together, these components form a closed-loop architecture that integrates cognitive delegation, speculative decoding, adaptive scheduling, and context compression—optimizing the reasoning lifecycle in its entirety.
@@ 2 Related Work
Large language models (LLMs) often form the reasoning core of agent systems, but their computational and memory demands limit interactive efficiency. Techniques such as model quantization, pruning, and low-rank adaptation have been widely investigated to reduce inference latency. Post-training quantization methods (e.g., W8A8, GPTQ, SmoothQuant) have achieved tangible speedups but may degrade reasoning accuracy on multi-step or symbolic tasks [dettmers2023qlora, frantar2023gptq]. Recent works like QLoRA [dettmers2023qlora] and AWQ [lin2024awq] aim to maintain performance via adaptive weight grouping and normalization-aware tuning. In addition, distillation-based models [hinton2015distilling, wang2024tinyllm] compress high-capability agents into lightweight student models for low-latency scenarios. These methods focus primarily on per-token acceleration but do not directly address multi-turn agent bottlenecks such as context growth or reasoning instability.
@@
Context explosion is a major concern in agent execution loops, as the history of reasoning, tool calls, and observations grows over time. To handle this, retrieval-augmented generation (RAG) [lewis2020rag, gao2023retrieval] integrates external memory to avoid long prompt reuse. Hierarchical summarization and ephemeral memory [xu2024memorybank, chen2024retentive] further compress intermediate interactions while preserving key task-specific cues. However, summarized or lossy memories can introduce semantic drift between an agent’s reasoning state and its retrieved context, a phenomenon recently observed in multi-turn setups [park2023generativeagents]. More persistent memory mechanisms—such as vector-space long-term memory stores or semantic caches—are being explored to ensure contextual continuity across sessions [yuan2024longcontext, liu2024memgpt].
@@
Reasoning efficiency can be improved by both algorithmic and architectural advances. Tree-of-Thoughts (ToT) [yao2024treeofthoughts] and Reflection Loop agents [shinn2024reflexion] structure inference into tree or iterative reasoning loops, trading additional calls for stability and correctness. To mitigate the resulting latency, hybrid strategies such as speculative decoding [leviathan2023fast, sun2025scalingupspeedingup], early-exit reasoning [zhang2024eegpt], and dynamic reasoning depth [xie2024dynapipe] have been introduced. These frameworks parallelize or truncate reasoning trajectories while keeping semantic fidelity, yielding near-optimal balance between compute and accuracy.
@@
At the system level, multi-agent orchestration frameworks [schick2023toolformer, gao2024autogen] expose severe heterogeneity in request length and tool usage patterns. Efficient scheduling algorithms—such as cache-aware dispatching [huang2024kvprefetch], hybrid batching [li2024llmbatching], or pipelined execution [raja2024deepinfra]—have shown significant throughput gains for heterogeneous workloads. Moreover, persistent KV-cache management [dao2024flashdecoding, li2025kvtuner] and prefix-sharing mechanisms improve utilization for long interactions, mitigating recomputation costs during prefill stages. Recent open systems like vLLM [kumar2024vllm] and SGLang [liu2024sglang] provide foundational infrastructures supporting efficient agent orchestration across distributed environments.
@@ 3 Agent Workflows
In certain multi-agent frameworks featuring explicit agent orchestration (e.g., DeepDiverV2 [openpangu_deepdiver_v2]), the central controller is referred to as the Planner Agent. It is responsible for decomposing complex objectives into well-defined sub-tasks and coordinating the execution flow among subordinate agents. Once the overall task plan is established, the Planner Agent dynamically instantiates and supervises specialized sub-agents to perform targeted operations—such as the Information Seeker Agent, which conducts knowledge retrieval and evidence collection, and the Writer Agent, which integrates retrieved information, synthesizes insights, and produces the final structured output. This hierarchical orchestration design enables collaborative reasoning and division of labor among heterogeneous agents, thereby improving both the efficiency and robustness of the overall problem-solving process.
@@ 5.4.1 Methodology
In agent inference, requests derived from the same user query—as well as those from different users with similar intents—exhibit strong correlations among the planner agent, executor agents, and tool agents. In deep research settings, retrieved search content is frequently revisited during the model’s reasoning process, making such environments well-suited for lightweight n-gram–based speculative decoding algorithms. We employ suffix automaton (SAM) [hu2024samdecodingspeculativedecoding] as a lightweight, hot-swappable speculative mechanism designed to provide lossless acceleration for small-batch and long-tail tasks.
@@ 6.1 Experimental Setup
  • Tasks: Experiments were conducted on the BrowseComp-zh (Chinese version) [zhou2025browsecomp] benchmark, which involves multi-step tasks requiring web searches.
  • Models: The primary model used was openPangu-DeepDiverV2-7B [openpangu_deepdiver_v2], with openPangu-7B [openpangu-embedded-7b] used for context compression in AgentCompress and openPangu-DeepDiverV2-38B [openpangu_deepdiver_v2] used for the large model in AgentCollab.
  • Infrastructure: vLLM-Ascend v0.9.1 [vllm-ascend] on Ascend 910B3 NPUs.

### v2

@@ 1 Introduction
Autonomous agents [1, 8], powered by large language models (LLMs), are poised to revolutionize domains ranging from deep scientific research and iterative software engineering to knowledge-intensive decision-making. Unlike conventional single-turn LLM applications—where the model generates a one-off response and terminates—autonomous agents operate within a persistent Think–Act–Observe loop. In this cycle, the agent formulates plans, interacts with external tools (e.g., search engines, simulators, code interpreters), observes outcomes, and iteratively reasons based on accumulated context. While this continuous, reflection-driven process endows agents with remarkable problem-solving flexibility, it imposes severe latency penalties and computational overheads. Each “Think” step typically triggers a large model inference involving extended contexts, parallel requests, and complex reasoning paths, resulting in cumulative delays that hinder real-time deployment and scalability.
A fundamental dichotomy exists between agent acceleration and traditional LLM inference optimization. Conventional speedups—such as quantization [14], pruning [1, 22, 21], or speculative decoding [27]—predominantly focus on intra-step efficiency, aiming to maximize tokens per second (TPS). In contrast, the efficiency of an autonomous agent is defined by the inter-step time-to-completion of multi-turn tasks. A marginal gain in decoding speed is negligible if the agent, due to degraded reasoning quality, requires dozens of additional correction cycles. Thus, optimizing an agent system demands a paradigm shift: bridging intra-step acceleration (faster inference) with inter-step efficacy (fewer steps to solution) [34, 16].
Consequently, traditional metrics like TPS can be perilously misleading in agentic contexts. An agent is not merely a passive text generator but an active reasoning system where progress is path-dependent [9]. A “faster but less accurate” model often regresses overall performance by introducing factual errors, invalid tool calls, or incoherent summaries that cascade through subsequent turns. Such errors necessitate retries and redundant inference, negating any low-level speedups. Therefore, agent performance must be evaluated through holistic metrics that consider End-to-End Task Latency (wall-clock time), Total Computational Cost (aggregate FLOPs), and Turn Efficiency (reasoning stability).
@@
We argue that accelerating agent inference requires a synergistic, multi-layered framework spanning both the reasoning process and the inference engine. To this end, we introduce AgentInfer, a hierarchical framework that systematically enhances agent efficiency. While our principles apply to general agentic systems (e.g., coding or dialogue agents), we validate them within the demanding context of Deep Research Agents. At the reasoning level, we propose AgentCollab, a self-evaluation-driven dual-model mechanism. It delegates routine inference and tool execution to a smaller model, while a larger model handles strategic planning and rescues stalled trajectories. This is governed by structured Progress Check signals, which trigger escalation only when necessary, thereby conserving expensive capacity. Complementarily, AgentCompress implements lightweight context summarization, pruning redundant reasoning traces to maintain compact yet semantically complete histories. At the inference-engine level, AgentSAM advances speculative decoding [7] by using a suffix automaton to efficiently reuse token-level predictions across agent sessions. Simultaneously, AgentSched introduces a hybrid, KVCache-aware scheduling policy. By utilizing a dynamic shadow-price controller, it adaptively switches between Shortest-Job-First and cache-preserving modes, balancing latency and cache hit rates under dynamic workloads. Together, these components form a closed-loop architecture that integrates cognitive delegation, speculative decoding, adaptive scheduling, and context compression—optimizing the reasoning lifecycle in its entirety.
@@ 2 Related Work
Large language models (LLMs) often form the reasoning core of agent systems, but their computational and memory demands limit interactive efficiency. Techniques such as model quantization, pruning, and low-rank adaptation have been widely investigated to reduce inference latency. Post-training quantization methods (e.g., W8A8, GPTQ, SmoothQuant) have achieved tangible speedups but may degrade reasoning accuracy on multi-step or symbolic tasks [3, 4]. Recent works like QLoRA [3] and AWQ [15] aim to maintain performance via adaptive weight grouping and normalization-aware tuning. In addition, distillation-based models [6, 31] compress high-capability agents into lightweight student models for low-latency scenarios. These methods focus primarily on per-token acceleration but do not directly address multi-turn agent bottlenecks such as context growth or reasoning instability.
@@
Context explosion is a major concern in agent execution loops, as the history of reasoning, tool calls, and observations grows over time. To handle this, retrieval-augmented generation (RAG) [13, 10] integrates external memory to avoid long prompt reuse. Hierarchical summarization and ephemeral memory [39, 2] further compress intermediate interactions while preserving key task-specific cues. However, summarized or lossy memories can introduce semantic drift between an agent’s reasoning state and its retrieved context, a phenomenon recently observed in multi-turn setups [20]. More persistent memory mechanisms—such as vector-space long-term memory stores or semantic caches—are being explored to ensure contextual continuity across sessions [23, 32].
@@
Reasoning efficiency can be improved by both algorithmic and architectural advances. Tree-of-Thoughts (ToT) [36] and Reflection Loop agents [25] structure inference into tree or iterative reasoning loops, trading additional calls for stability and correctness. To mitigate the resulting latency, hybrid strategies such as speculative decoding [12, 26], early-exit reasoning [28], and dynamic reasoning depth [33] have been introduced. These frameworks parallelize or truncate reasoning trajectories while keeping semantic fidelity, yielding near-optimal balance between compute and accuracy.
@@
At the system level, multi-agent orchestration frameworks [24, 30] expose severe heterogeneity in request length and tool usage patterns. Efficient scheduling algorithms—such as cache-aware dispatching [17], hybrid batching [35], or pipelined execution [37]—have shown significant throughput gains for heterogeneous workloads. Moreover, persistent KV-cache management [5, 14] and prefix-sharing mechanisms improve utilization for long interactions, mitigating recomputation costs during prefill stages. Recent open systems like vLLM [11] and SGLang [38] provide foundational infrastructures supporting efficient agent orchestration across distributed environments.
@@ 3 Agent Workflows
In certain multi-agent frameworks featuring explicit agent orchestration (e.g., DeepDiverV2 [18]), the central controller is referred to as the Planner Agent. It is responsible for decomposing complex objectives into well-defined sub-tasks and coordinating the execution flow among subordinate agents. Once the overall task plan is established, the Planner Agent dynamically instantiates and supervises specialized sub-agents to perform targeted operations—such as the Information Seeker Agent, which conducts knowledge retrieval and evidence collection, and the Writer Agent, which integrates retrieved information, synthesizes insights, and produces the final structured output. This hierarchical orchestration design enables collaborative reasoning and division of labor among heterogeneous agents, thereby improving both the efficiency and robustness of the overall problem-solving process.
@@ 5.4.1 Methodology
In agent inference, requests derived from the same user query—as well as those from different users with similar intents—exhibit strong correlations among the planner agent, executor agents, and tool agents. In deep research settings, retrieved search content is frequently revisited during the model’s reasoning process, making such environments well-suited for lightweight n-gram–based speculative decoding algorithms. We employ suffix automaton (SAM) [7] as a lightweight, hot-swappable speculative mechanism designed to provide lossless acceleration for small-batch and long-tail tasks.
@@ 6.1 Experimental Setup
  • Tasks: Experiments were conducted on the BrowseComp-zh (Chinese version) [40] benchmark, which involves multi-step tasks requiring web searches.
  • Models: The primary model used was openPangu-DeepDiverV2-7B [18], with openPangu-7B [19] used for context compression in AgentCompress and openPangu-DeepDiverV2-38B [18] used for the large model in AgentCollab.
  • Infrastructure: vLLM-Ascend v0.9.1 [29] on Ascend 910B3 NPUs.